Find the compound interest if RM1000 is invested for five years at an annual interest rate of 7%
compounded :


a) quarterly
b) monthly
c) weekly
Answer #1

quarterly

i) principal = $ 1000

rate = 7%

time = 5 years

A = 1000 ( 1 + .07/4)^20

A = $ 1414.78

ii) effective rate = ( 1 + .07/4 )^4 - 1

= 7.186 %

b) monthly

A = 1000 ( 1 + .07/12 )^60

A = $ 1417.63

ii) effective rate = ( 1+ .07/12 )^12 - 1

= 7.229 %

c) weekly

A = 1000 ( 1 + .07/52 )^52*5

A = $ 1418.73

ii) effective rate = ( 1 + .07/52 )^52 - 1

= 7.246%

d) daily

A = 1000 ( 1 + .07/365 )^365*5

A = $ 1419.02

ii) effective rate = ( 1 + .07/365)^365 - 1

= 7.25 %


answered by: Opicyc